OpenStudy (anonymous):

Write the polynomial in factored form: X^3 + 7x^2 + 15x + 9

OpenStudy (anonymous):

replace x by -1 and see what you get

OpenStudy (anonymous):

If you get zero, this will imply that x+1 is a factor

OpenStudy (anonymous):

You can divide by x+1 and obatain a quadratic that might be easier to factor

OpenStudy (anonymous):

How do I do that? I'm confused because it's not a trinomial

OpenStudy (jhannybean):

Ahh, this relates to the Rational Root Theorem, I believe, yes?

OpenStudy (anonymous):

If you divide your polynomial by x +1, you get \[ \frac{x^3+7 x^2+15 x+9}{x+1}=x^2+6 x+9 \]

OpenStudy (anonymous):

Can you factor \[ x^2+6 x+9 \]

OpenStudy (anonymous):

Yes it does

OpenStudy (anonymous):

Notice that \[ x^2+6 x+9=(x+3)^2 \]

OpenStudy (anonymous):

I know how to factor that yea, I don't really get how you divide by x-1

OpenStudy (anonymous):

So finally \[ x^3+7 x^2+15 x+9=(x+1) (x+3)^2 \]

OpenStudy (anonymous):

Use synthetic division

OpenStudy (jhannybean):

The way @eliassaab found -1 , @yankeeez , is by taking all the possible factors of the function where \( x= \pm \frac{p}{q}\) where p is the constant infront of \(x^3\) and \(q = 9\). This gives -1 as a possible candidate for factoring out the polynomial in preparation for reduction.

OpenStudy (anonymous):

http://mathworld.wolfram.com/SyntheticDivision.html